the total volume of the gas bags of the german girigible hindenberg, was 2.0 x 10^5 m^3. How many grams of H2 would be required to fill them at 20.0C and 1.00 atm pressure?
Answer #1

Given:
P = 1.0 atm
V = 200000 m^3
= (200000*1000) L
= 200000000 L
T = 20.0 oC
= (20.0+273) K
= 293 K

find number of moles using:
P * V = n*R*T
1 atm * 200000000 L = n * 0.08206 atm.L/mol.K * 293 K
n = 8.318*10^6 mol

Molar mass of H2 = 2.016 g/mol

use:
mass of H2,
m = number of mol * molar mass
= 8.318*10^6 mol * 2.016 g/mol
= 1.677*10^7 g
Answer: 1.68*10^7 g